OpenStudy (anonymous):

This question is about functions f(x) = x^2 - 13x - 30 and g(x) = -3x^3 + 7x. Calculate f(x) times g(x). Simplify the answer as much as possible.

myininaya (myininaya):

\[f\cdot g=(x^2-13x-30)(-3x^3+7x)=x^2(-3x^3+7x)-13x(-3x^3+7x)-30(-3x^3+7x)\]

myininaya (myininaya):

now apply the distributive property for each of the three terms

myininaya (myininaya):

and then combine like terms

OpenStudy (anonymous):

a little long there myininaya

myininaya (myininaya):

and woolah!

jimthompson5910 (jim_thompson5910):

\[\large (f(x))(g(x))\] \[\large (x^2 - 13x - 30)(-3x^3 + 7x)\] \[\large (-3x^3 + 7x)(x^2 - 13x - 30)\] \[\large -3x^3(x^2 - 13x - 30)+7x(x^2 - 13x - 30)\] \[\large -3x^5 +39x^4+ 90x^3+7x^3 -91x^2 -210x\] \[\large -3x^5 +39x^4+ 97x^3 -91x^2 -210x\] So \[\large (x^2 - 13x - 30)(-3x^3 + 7x)=-3x^5 +39x^4+ 97x^3 -91x^2 -210x\]

myininaya (myininaya):

is it cut off? it doesn't show me any cutoffness

OpenStudy (anonymous):

maybe it is my browser. i see as far as -13x

myininaya (myininaya):

:(

myininaya (myininaya):

ok well just look at what jim has anything he ever does is perfect

jimthompson5910 (jim_thompson5910):

I find it better to expand (A+B+C)(D+E) to D(A+B+C)+E(A+B+C) instead of A(D+E)+B(D+E)+C(D+E) since it's shorter but you'll get the same answer either way